-6p - 21 = 3p - 12
Found 2 solutions by stanbon, Maths68:
Answer by stanbon(75887)

-6p - 21 = 3p - 12
---
Add 6p to both sides
Add 12 to both sides
---
9p = -9
p = -1
========
Cheers,
Stan H.
===============

Answer by Maths68(1474)

-6p - 21 = 3p - 12
Add 21 both sides of the above equation
-6p - 21+21 = 3p - 12 + 21
Cancel out 21 from left hand side
-6p = 3p + 9
Subtract 3p from both sides of the above equation
-6p -3p = 3p - 3p + 9
Cancel out 3p from right hand side
-9p =9
Divide by -9 both sides of the above equation
-9p/-9 =9/-9
Cancel out 9 from both sides
p=-1
Check
Put the value of p in original equation
-6p - 21 = 3p - 12
-6(-1) - 21 = 3(-1) - 12
6-21=-3-12
-15=-15